OpenStudy (anonymous):

Defining e by: \[\int\limits_{1}^{e}{1\over x} dx = 1\] show that \[\int\limits_{1}^{t}{1\over x} dx = ln t\]

OpenStudy (jamesj):

In the second integral, we want to change variables such that the limits of integration are no longer 1 to t but 1 to e and we have the same integrand. That being the case, think about what substitution/change of variable works.

OpenStudy (anonymous):

\[x= {tu \over e}\] when x = t, u =e, but when x =1 u isn't 1

OpenStudy (jamesj):

No, it's not. This is a bit tricky. Think of something exponential. Let u be the new variable. Then with u = f(x) we want f(1) = u = 1 f(t) = u = e

OpenStudy (jamesj):

or the other way around x = g(u) where g(1) = x = 1 g(e) = x = t

OpenStudy (jamesj):

hint: \[ t = e^{\ln t} \]

OpenStudy (jamesj):

and 1 = 1^(ln t)

OpenStudy (anonymous):

so \[x = e ^{\ln u }\]

OpenStudy (jamesj):

No, because when x = t, you have u = e^t

OpenStudy (anonymous):

no thats not right

OpenStudy (jamesj):

Again, we want g where x= g(u) 1 = x = g(1). t = x = g(e) Hint: e^(ln t) = t

OpenStudy (jamesj):

1 = x = g(1). Hint: 1^(ln t) = 1 t = x = g(e) Hint: e^(ln t) = t

OpenStudy (anonymous):

i'm going to sleep on it, I understand what i need to get to ....

OpenStudy (jamesj):

ok.

OpenStudy (anonymous):

thanks for your help

OpenStudy (anonymous):

u = e ^ ln x

OpenStudy (jamesj):

No. It will be easier if you try and write as I indicate above x = g(u)

OpenStudy (jamesj):

1 = x = g(1). Hint: 1^(ln t) = 1 t = x = g(e) Hint: e^(ln t) = t And remember that t is a constant.

OpenStudy (anonymous):

thanks once again, brain is drained!

OpenStudy (jamesj):

ok. I admire your discipline in not asking me the answer ... yet!

OpenStudy (anonymous):

i'm not going to!

OpenStudy (anonymous):

how about x = u ^ ln t so when x = 1 u =1, when x = t, u = e

OpenStudy (jamesj):

Looks promising. Now, does it work? What is dx and hence the integral \[ \int_1^t \frac{dx}{x} \]

OpenStudy (anonymous):

\[\ln t \int\limits_{1}^{e} {1 \over u} du\]

OpenStudy (anonymous):

which gives the answer

OpenStudy (anonymous):

blimey that was a hard substitution for A level

OpenStudy (jamesj):

Yes. Good.

OpenStudy (anonymous):

or maybe i'm just thick

OpenStudy (jamesj):

This was a curious question. It it quite frequent to define ln t as that integral. But I've never seen anyone asked to derive this from a different starting point.

OpenStudy (anonymous):

that was fun working it out, well with your help of course

OpenStudy (jamesj):

I'm a mathematician and if it makes you feel any better, it took me a couple of minutes to find the substitution as well. Quite unusual.

OpenStudy (anonymous):

oh good! makes me feel a bit better

OpenStudy (anonymous):

hope yet

OpenStudy (anonymous):

right i can sleep easy now thanks James

OpenStudy (jamesj):

have a good weekend.

OpenStudy (anonymous):

and you